a chemist needs 10 liters of a 50% salt solution. all that is available is a 20% salt solution and a 70% salt solution. how much of each of the 2 solutions should she mix to obtain her desired solution?
a) 5 liters of the 20% solution; 5 liters of the 70% solution
b) 3 liters of the 20% solution; 7 liters of the 70% solution
c) 2 liters of the 20% solution; 8 liters of the 70% solution
d) 4 liters of the 20% solution; 6 liters of the 70% solution

Answer by Paul(988)

20x+70(10-x)=50(10)
20x+700-70x=500
-50x=-200
x=4
10-4=6
Hence, for 20% there should be 4L of solution added and for 70% there should be 6L if solution added